OpenStudy (anonymous):

Find a formula for the inverse of the following equation:

OpenStudy (anonymous):

\(\ \huge y=ln(x+3) \).

OpenStudy (anonymous):

Let f(x) = ln(x + 3) ff^-1 (x) = x ln(f^-1(x) + 3) = x f^-1(x) + 3 = e^x f^-1(x) = e^x - 3 The inverse of y is e^x - 3

OpenStudy (anonymous):

? Don't you just need to switch the x and y around, and then simplify?

OpenStudy (anonymous):

I don't really understand what you did...

OpenStudy (anonymous):

No, @seashell your previous answer was the correct answer, but I don't understand how you got to it

OpenStudy (anonymous):

i just understand it this weird way :/

OpenStudy (anonymous):

ln (x+3)=y just take the e^() eponent of both sides e^(ln (x+3))=e^y x+3=e^y left hand side turns to x+3 x=e^y-3 subtract from both sides 3 now you can swap variables y=e^x-3 Hope I got it right